OpenStudy (anonymous):

Use basic identities to simplify the expression. cos^2Θ/sin^2Θ + cscΘ sinΘ

OpenStudy (anonymous):

\[ \frac{cos^2\theta}{sin^2\theta} + csc\theta sin\theta = tan^2\theta + 1 = sec^2\theta\]

OpenStudy (anonymous):

Okay but how did you do it?

OpenStudy (anonymous):

How did you get cscΘ

OpenStudy (anonymous):

nv whoops haha

OpenStudy (anonymous):

i meant tan^2Θ+1

OpenStudy (anonymous):

by using the identity: \[ sec^2\theta - tan^2\theta = 1\]